# Lindblad-like quantum tomography for non-Markovian quantum dynamical maps

### v1

@@ II Learning time-local master equations by maximum-likelihood estimation
III Lindblad-like quantum tomography for non-Markovian dephasing
@@
Figure 1: Scheme for $\ell$QT protocol of a dynamical dephasing map: In the case of pure dephasing, the $\ell$QT scheme reduces to a problem of Ramsey estimation. In the QIP, the qubit is initialized in a single state in the equator of Bloch’s sphere $\rho_{0,0}=\outerproduct{+}{+}$ | by a resonant driving inducing a $\pi/2$ pulse on a reference state $\rho_{0}=\outerproduct{0}{0}$ |. The qubit then evolves for different times $t_{i}:i\in\mathbb{I}_{t}$, after which it is subjected to another resonant $\pi/2$ pulse, which effectively rotates to the Pauli measurement basis $x$. As a consequence of the dephasing noise, the $m_{x}\in\{0,1\}$ outcomes are collected as relative frequencies $\{f_{+,i,m_{x}}\}$ which will decay in time. In the lower half, we show the classical part of the estimation, which starts by solving the time-local master equation for a particular parametrized power spectral density $S_{\bm{\theta}}(\omega)$ and filter function. This leads to the theoretically predicted probabilities $\{p_{+,i,m_{x}}(\bm{\theta})\}$. These probabilities and the relative frequencies are fed into a log-likelihood cost function, which must be minimized in order to obtain the estimate. By calculating the covariance matrix, we can estimate the precision, and search for the optimal measurement times, which will depend on how many noise parameters we aim at estimating, and changes with their specific values. In a Bayesian approach, rather than minimizing a cost function built from the statistical likelihood, one updates the knowledge about the noise parameters sequentially by selecting the evolution times that are expected to provide most information at each Bayesian step.
@@
Let us formulate the $\ell$QT for the time-local master equation (8) for the dephasing of a single qubit, including temporal correlations. As discussed in Appendix A, either in a semi-classical or quantum-mechanical model of pure dephasing, the qubit dynamics can be described by a time-local master equation (8) with a simple Hamiltonian $H(t)=\textstyle\frac{1}{2}\omega_{0}\sigma_{z}$ and a single jump operator $L(t)=\sigma_{z}$, both of which are time independent
@@
In a semi-classical model, $C(t,t^{\prime})=\mathbb{E}[\delta\tilde{\omega}(t)\delta\tilde{\omega}(t^{\prime})]$ is the auto-correlation function of a stochastic process $\delta\tilde{\omega}(t)$ representing frequency fluctuations. This master equation is the result of averaging over the stochastic process $\rho(t)=\mathbb{E}[\tilde{\rho}(t)]$, and is valid for a random process in a second-order cumulant expansion know as the fast-fluctuation expansion or, alternatively, for a Gaussian random process with arbitrary correlation times $\tau_{\rm c}$, as discussed in App. A. Alternatively, for a fully quantum-mechanical dephasing model, $C(t,t^{\prime})=\textstyle\frac{1}{2}{\rm Tr}_{B}\{B_{I}(t)B_{I}(t^{\prime})\rho^{\rm ss}_{\rm B}\}$ is the auto-correlation function on the stationary state of the environment/bath $\rho^{\rm ss}_{\rm B}$, which induces fluctuations in the qubit frequency via the bath operators $B_{I}(t)$. In this case, the time-local master equation is the result of tracing over the bath degrees of freedom $\rho(t)={\rm Tr}_{B}\{\rho_{\rm SB}(t)\}$, and is valid in a second-order cumulant expansion also discussed in App. A. Assuming wide-sense stationarity, i.e., the mean of the stochastic process is constant and the auto-correlation function only depends on $t^{\prime}-t$, we can introduce the power spectral density (PSD) of the noise
@@
It will be useful to define the symmetrized auto-correlation function and the symmetrized PSD as $\bar{C}(t,t^{\prime})=\frac{1}{2}(C(t,t^{\prime})+C(t^{\prime},t))$ and $\bar{S}(\omega)=\frac{1}{2}(S(\omega)+S(-\omega))$, since for dephasing noise only the symmetric part of the auto-correlation function and the symmetric part of the PSD influence the time evolution, as discussed in App. A.
@@
In light of the $\ell$QT estimation problem of Eq. (11), we do not need to consider arbitrary time-dependent functions for the Hamiltonian and dissipation matrix, but can actually find an effective parametrization that reduces drastically the search space. Considering Eq. (15), the dephasing rate will depend on a certain number $n$ of real-valued noise parameters $\bm{\theta}_{\!\star}\in\Theta=\mathbb{R}^{n}$ via the PSD $S_{\bm{\theta}_{\!\star}}\!(\omega)$, where we have made explicit its parametrization. Since the Hamiltonian of the qubit is very simple, and only depends on the transition frequency $\omega_{0}$ that is known with a high precision using spectroscopic methods, it need not be included in the learning. Technically, this means that we can assume that the driving used to initialize and measure the qubit is resonant with the transition (see Fig. 1), and work in the rotating frame of App. A. Likewise, we have only one possible jump operator $L=\sigma_{z}$, such that the learning can focus directly on the estimation $\bm{\hat{\theta}}$ of the actual noise parameters $\bm{{\theta}}_{\star}$, which translate into the estimation of the dephasing rate $\hat{\gamma}(t)$ via Eq. (15). Therefore, $\ell$QT becomes a non-linear minimization problem for the cost function (9). Before giving more details on this problem, we discuss relevant properties of the dephasing quantum dynamical map.
@@
For a white-noise noise model with a vanishing correlation time $\tau_{\rm c}=0$, one has a flat PSD $S(\omega)=c$ and a constant dephasing rate $\gamma(t)=c/2$. This leads to an exponential decay of the coherences $p_{+,i,x}(m_{x})=\textstyle\frac{1}{2}\big{(}1+(-1)^{m_{x}}\mathord{\rm e}^{-t_{i}/T_{2}}\big{)}$, where $m_{x}\in\{0,1\}$ correspond to the projective measurements on $\ket{+},\ket{-}$, respectively, and we have introduced a decoherence time $T_{2}=2/c$. If the real system is affected by white dephasing noise, there is thus a single noise parameter to learn $\theta_{\star}=c_{\star}$ or, alternatively, the real decoherence time $\theta_{\star}=T_{2\star}$. We note that this procedure is in complete agreement with the LQT based on the corresponding Lindblad master equation (2). On the other hand, for a time-correlated dephasing noise with a structured PSD, the coherence decay will generally differ from the above exponential law, with the exception of the long-time regime $t_{i}\gg\tau_{\rm c}$, where $p_{+,i,x}(m_{x})\approx\textstyle\frac{1}{2}\big{(}1+(-1)^{m_{x}}\mathord{\rm e}^{-t_{i}/T_{2}}\big{)}$ and one finds an effective decoherence time controlled by the static part of the PSD $T_{2}=2/S(0)$. As $t_{i}$ increases towards $\tau_{\rm c}$, the decay will no longer be a time-homogeneous exponential, which can actually be a consequence of (but not a prerequisite for) a non-Markovian quantum evolution. In this more general situation, we will have more noise parameters $\bm{\theta}_{\star}$ to learn.
@@
is non-Markovian when, at some intermediate time, there is no CPTP map $\mathcal{E}^{\rm TL}_{t,t^{\prime}}$ such that $\mathcal{E}_{t,0}^{\rm TL}=\mathcal{E}^{\rm TL}_{t,t^{\prime}}\circ\mathcal{E}^{\rm TL}_{t^{\prime},0}$ (7). In the above expression (22), we have introduced the following time-dependent probability for the occurrence of phase-flip errors

### v2

@@ Results
As stated in Eq. (7) we can follow a frequentist approach similarly to the one in Markovian Lindblad quantum tomography [87, 89, 90], but now taking into account the time-dependence of the decay rates, which will require an a priori selection of the evolution times at which the system is probed. In light of the $\ell$QT estimation problem of Eq. (7), we do not need to consider arbitrary time-dependent functions for the Hamiltonian and dissipation matrix, but can actually find an effective parametrization that reduces drastically the search space. This means that the minimization in Eq. (7) can be over the parameters $\bm{\theta}$ of the Hamiltonian and dissipation matrix. Although several intermediate times can be taken, a single time instant actually suffices for an accurate learning in LQT [90], which makes it simpler and computationally lighter. As the noise becomes time correlated, reconstructing the more complex time-dependent dynamics via $\ell$QT, demands multiple time steps to achieve high accuracies. LQT has in principle the same number of parameters $d^{2}(d^{2}-1)$ to be learnt as quantum process tomography, and to do so requires at least $d^{2}(d^{2}-1)$ measurement configurations. For the general case, where arbitrary time-dependent rates need to be estimated, a naive blind-search approach would require repeated LQT at many time points, leading to high sample complexity. In contrast, $\ell$QT’s use of a microscopic parameterization reduces the need for extensive measurements by imposing physical constraints, which partially addresses the sample complexity challenge and makes it a more scalable method. Thus in parameterized $\ell$QT we will need at least as many measurement configurations as the number of parameters to estimate.
Alternatively to this frequentist approach, we can also consider a Bayesian approach that exploits a physically-motivated prior knowledge about these noise parameters, which can be represented by a certain probability distribution. After performing measurements on the system at certain instants of time and certain measurement basis, we update this probability distribution with the new acquired information, a process that is repeated until reaching a target accuracy for the parameter estimation. The Bayesian approach has the advantage of choosing, at each step, the most convenient subsequent time and measurement basis at which to measure by maximizing the information one would gain. In principle, this can lead to a reduction in the number of measurements required to reach a certain accuracy with respect to those required by the frequentist approach. In practice, however, the non-Markovianity of the quantum evolution can modify this argument, as our probabilistic account of the model parameters can be affected by the actual time correlations of the underlying random process.
Frequentist approach to non-Markovian inference
In this section, we focus on the frequentist approach, which builds on the relative frequencies of observed outcomes. The $\ell$QT cost function of Eq. (5) can be rewritten as follows
$\mathsf{C}_{\mathrm{TL}}(\bm{\theta})=-\sum_{s,i,b}N_{s,i,b}\,\sum_{m_{b}}\,\tilde{f}_{s,i,b}(m_{b}\,|\bm{\theta}_{\star})\log p^{\rm TL}_{s,i,b}(m_{b}|\bm{\theta}),$ (8)
where $\tilde{f}_{s,i,b}(m_{b}\,|\bm{\theta}_{\star})=N_{s,i,b,m_{b}}/N_{s,i,b}$ is the ratio of the number of $m_{b}$ measurement outcomes observed $N_{s,i,b,m_{b}}$ to the total of $N_{s,i,b}$ shots collected at the instant of time $i$ and initial state $s$ when measuring the system with the POVM element $b$. Our notation remarks that these relative frequencies carry information about the real noise parameters $\bm{\theta}_{\!\star}$ we aim at estimating. In addition, the estimator depends on $p_{s,i,b}(m_{b}|\bm{\theta})$ shown in Eq. (6), where we make explicit the dependence on the parametrized noise.
The minimization in Eq. (7) can explicitely written as a minimization over the noise parameters. As a consequence, the frequentist approach can be recast as a statistical problem of parameter point estimation [104], namely
$\bm{\hat{\theta}}_{\rm F}=\texttt{argmin}_{\bm{\theta}}\{\mathsf{C}_{\mathrm{TL}}(\bm{\theta}):\,\bm{\theta}\in\Theta=\mathbb{R}^{n}\}.$ (9)
Instead of the general $\ell$QT learning over $d^{2}(d^{2}-1)$ parameters, which increase exponentially $d=2^{n_{q}}$ with the number of qubits $n_{q}$ and can arbitrarily change in time, our procedure revolves around the estimation of $n$ noise parameters, which are independent of the system size and the evolution times. On the other hand, the imprecision of our estimates will indeed depend on our choice of the evolution times, forcing us to go beyond the LQT single-time estimator [90], and actually measure at optimal times, initial states and measurement elements for which the estimation imprecision can be minimized. Let us note that the conditions to minimise this cost function are the same as those that minimize the Kullback-Leibler divergence [106, 107], which is the following relative entropy
between the experimental $p_{1,k}\in\{\tilde{f}_{s,i,b}(m_{b}\,|\bm{\theta}_{\star})\}$ and parameterized theoretical $p_{2,k}\in\{p^{\rm TL}_{s,i,b}(m_{b}\,|\bm{\theta})\}$ probability distributions, provided one considers variations with respect to the estimation parameters $\bm{\theta}$.
We note that the expectation values are taken with respect to the probability distributions for the measurements of $p_{s,i,b}(m_{b})$ which, implicitly, also have the stochastic average over the random dephasing noise in the semi-classical model, or a partial trace over the environment in the quantum-mechanical one. The nice property of the maximum-likelihood estimator is that it is asymptotically unbiased, such that $\bm{\mathsf{B}}_{\bm{\theta}_{\star}}\!\!(\bm{\theta}_{\rm F})=\mathbb{E}[\bm{\hat{\theta}}_{\rm F}]-\bm{\theta}_{\star}\to\bm{0}$ for a sufficiently large $N_{\rm shot}$. Moreover, its asymptotic covariance matrix saturates the Cramér-Rao bound [108] relating the estimation precision to the Fisher information matrix, which quantifies the amount of information in $\mathbb{D}$ about the unknown parameters. If we momentarily assume that the measurements occur at a single instant of time $t=t_{i}$, a single initial state and a single measurement basis with outcomes that are independent and identically distributed, the covariance matrix becomes $\mathrm{Cov}(\bm{\hat{\theta}}_{\rm F})\approx(N_{s,i,b}{I}_{F,s,i,b}(\bm{\theta}_{\!\star}))^{-1}$, where
$[{I}_{F,s,i,b}(\bm{\theta}_{\!\star})]_{n_{1},n_{2}}\!=\mathbb{E}\!\!\left[\left.\!\frac{\partial{\rm log}\big{(}p_{s,i,b}^{\rm TL}(m_{b}|\bm{\theta})\big{)}}{\partial{\theta_{n_{1}}}}\!\frac{\partial{\rm log}\big{(}p_{s,i,b}^{\rm TL}(m_{b}|\bm{\theta})\big{)}}{\!\partial\theta_{n_{2}}}\!\right|_{\bm{\theta}_{\!\star}}\right]\!.$ (12)
In this work, we deal with the more general case in which we measure at several times, therefore the random variables are not identically distributed, and the total number of shots need not be the same for different measurement configurations, i.e., instants of time, initial states and measurement basis. In this case, we must take a linear combination of the Fisher information matrices of each measurement configuration weighted by the proportion of measurements taken at each one of these configurations  [109]. We thus obtain the asymptotic covariance matrix
$\mathrm{Cov}(\bm{\hat{\theta}}_{\rm F})\approx\Sigma_{\bm{\hat{\theta}}}\equiv\left[N_{\rm shot}\sum_{s,i,b}\frac{N_{s,i,b}}{N_{\rm shot}}{I}_{F,s,i,b}(\bm{\theta}_{\!\star})\right]^{-1},$ (13)
such that, the more the Ramsey estimator varies under changes of the noise parameters, the bigger the amplification of the noise parameter is and, thus, the smaller the imprecision one can achieve. As we can see, the imprecision of the estimate will scale with $1/\sqrt{N_{\rm shot}}$, such that the Ramsey estimator is asymptotically consistent in the $N_{\rm shot}\rightarrow\infty$ asymptotic limit [104]. The asymptotic statistics of the maximum-likelihood estimator is further explained in the Methods section.
We also note that, in this limit, the observed relative frequencies will be normally distributed, such that one can consider minimizing a weighted least-squares cost function. For instance, assuming a binary case where $m_{b}\in\{0,1\}$, corresponding to a qubit measurement, we have
$\mathsf{C}_{\mathrm{TL}}(\bm{\theta})\approx\sum_{s,i,b}\frac{\big{(}\tilde{f}_{s,i,b}(0\,|\bm{\theta}_{\star})-p_{s,i,b}^{\rm TL}(0|\bm{\theta})\big{)}^{2}}{\tilde{\sigma}^{2}_{f_{s,i,b}}},$ (14)
where $\tilde{f}_{s,i,b}(0\,|\bm{\theta}_{\star})=N_{s,i,b,0}/N_{s,i,b}$ is the ratio of the number of outcomes observed $N_{s,i,b,0}$ to the total of shots $N_{s,i,b}$ collected at initial state $s$, the instant of time $t_{i}$ and measurement element $b$. Here, $\tilde{\sigma}^{2}_{f_{s,i,b}}$ is the variance of these measured samples. These $\tilde{f}_{s,i,b}(0\,|\bm{\theta}_{\star})$ follow a binomial proportion distribution with mean given by $p_{s,i,b}^{\rm TL}(0|\bm{\theta_{\star}})$ and, in the limit of a large $N_{s,i,b}$, they can be approximated by a normal distribution with mean $p_{s,i,b}^{\rm TL}(0|\bm{\theta_{\star}})$ and variance $\sigma^{2}_{f_{s,i,b}}$ according to the central limit theorem. The expected variance of the measurement configuration $(s,i,b)$ is
$\sigma^{2}_{f_{s,i,b}}=\frac{p_{s,i,b}^{\rm TL}(0\,|\bm{\theta})\big{(}1-p_{s,i,b}^{\rm TL}(0\,|\bm{\theta})\big{)}}{N_{s,i,b}},$ (15)
since we are sampling from a binomial distribution. This approximation allows us to use a simpler weighted least-squares algorithm, such as the trust-region reflective algorithm implemented in SciPy, where we can optionally set some bounds for the parameters to be estimated. The computational complexity of the trust-region reflective algorithm is dominated in each iteration by solving the trust-region subproblem, which is typically $O(n^{3})$, with $n$ the number of parameters. The number of iterations depends on the complexity of the cost function, the accuracy of the initial guess, and the desired level of convergence. However, once the algorithm gets close to the solution, the remaining convergence is rapid, showing a quadratic rate, meaning the distance to the solution decreases at a rate proportional to the square of the distance at the previous step.
Once the properties of the Ramsey frequentist estimate $\bm{\hat{\theta}}_{\rm F}$ have been discussed, we can search for the optimal measurement times, initial states and measurement elements that would lead to a Ramsey estimator with the lowest possible imprecision for a given finite $N_{\rm shot}$. Depending on which parameter we are interested in, we may be interested in minimizing a particular component of the covariance matrix $[\Sigma_{\bm{\hat{\theta}}}]_{n_{1},n_{1}}$ in Eq. (13) or, alternatively, minimize its determinant as a whole. In the asymptotic limit in which $\bm{\hat{\theta}}_{\rm F}$ follows a multivariate normal distribution, $\det\Sigma_{\bm{\hat{\theta}}}$ is proportional to the volume enclosed by the covariance elliptical region, so it is a good measure of the dispersion of the distribution, and a good way to quantify the accuracy of the estimation. Sometimes, we will also use $\det\Sigma_{\bm{\hat{\theta}}}^{1/2n}$ instead, which can be more easily compared to the individual standard deviations $[\Sigma_{\bm{\hat{\theta}}}]_{n_{1},n_{1}}^{1/2}$, as both quantities scale with $1/\sqrt{N_{\rm shot}}$. Minimizing $\det\Sigma_{\bm{\hat{\theta}}}$ has also the advantage that the optimal measurement configurations obtained are independent of the parameters we want to determine, assuming the different parametrizations have the same number of parameters and that a coordinate transformation exists between parametrizations. In this case, the determinants of the covariance matrices are related by $\det\Sigma_{\bm{\hat{\theta}}}=\det\Sigma_{\bm{\hat{\theta}}^{\prime}}\det J^{2}$, with $J$ the Jacobian of the coordinate transformation between both parametrizations, which does not depend on the measurement configurations and therefore will have no influence in the minimization. Before presenting these results, we discuss an approach based on Bayesian inference [104].
Bayesian approach to non-Markovian inference
Rather than considering the relative frequencies as approximations of the underlying probability distribution with a certain fixed value of $\bm{\theta}_{\star}$, the idea of Bayesian inference is to quantify statistically our knowledge about the noise parameters, and how this knowledge gets updated as we collect more information via measurements. Hence, the noise parameters become continuous stochastic variables themselves $\bm{\theta}_{\star}\mapsto\bm{\tilde{\theta}}$ that take values $\bm{\theta}\in\Theta$ according to a prior probability density function (PDF) $\pi_{0}(\bm{\theta})$. This probability distribution quantifies our uncertainty about the noise parameters before making any measurement $\mathbb{D}_{0}=\emptyset$. At each $\ell>0$ Bayesian step, we measure the system enlarging the data set sequentially $\mathbb{D}_{\ell-1}\mapsto\mathbb{D}_{\ell}=\mathbb{D}_{\ell-1}\cup\delta\mathbb{D}_{\ell}$, where $\delta\mathbb{D}_{\ell}\subset\mathbb{D}=\{N_{i,s,b,m_{b}}\}$ contains a number of measurement outcomes $|\delta N_{\ell}|$ that is a fraction of the total $N_{\rm shot}$. These outcomes will be labeled as $\delta\mathbb{D}_{\ell}=\{N_{i_{\ell},s_{\ell},b_{\ell},m_{b_{\ell}}}\}$. The measurements in this data set are again binary Bernoulli trials, and can be described by a joint distribution of independent binomials $p_{\tilde{N}}(\delta\mathbb{D}_{\ell})$, defined in analogy to the likelihood function, but only extended to the configurations measured in the particular Bayesian step. The prior $(\ell-1)$-th probability distribution is then updated by using Bayes’ rule based on the parametrized probability distributions $p^{\rm TL}_{\ell}(\delta{\mathbb{D}}_{\ell}|\bm{\theta})$ being understood as probabilities conditioned on our statistical knowledge of the noise parameters
[…]
$\mathsf{R}_{\rm TL}(\hat{\tau})=\!\sum_{i_{\ell},s_{\ell},b_{\ell},m_{b_{\ell}}\in\delta\mathbb{D}_{\ell}}\int_{\Theta}\!{\rm d}^{n}\theta\,\pi_{\ell}(\bm{\theta})p^{\rm TL}_{\ell}\big{(}\delta\mathbb{D}_{\ell}|\bm{\theta}\big{)}(\hat{\tau}-\bm{\theta})^{2}.$ (18)
In addition to the expectation value in Eq. (17), since we have the updated probability distributions, we can quantify how our uncertainty about the noise parameters changes via the associated covariance matrix $[\mathrm{Cov}(\bm{\hat{\theta}}_{{\rm B}})]_{n_{1},n_{2}}$ or any other statistics, regardless of the size of the Bayesian data set $\mathbb{D}_{\ell}$. This differs from our previous arguments for covariance of the maximum likelihood estimate $\bm{\hat{\theta}}_{\rm F}$ in Eq. (13), which require working in the asymptotic regime $N_{\rm shot}\rightarrow\infty$. In experimental situations in which this regime cannot be reached, we note that one could use Monte Carlo sampling techniques to estimate precision of $\bm{\hat{\theta}}_{\rm F}$ [110, 111, 77], although these deal with the estimator based on the full likelihood function.
Another crucial difference of the Bayesian approach is that, instead of choosing a predefined set of evolution times, initial states and measurement basis, we can find the optimal time at which we should measure to maximize the information gain at each Bayesian step. For each update, we thus solve for
$(i_{\ell},s_{\ell},b_{\ell})=\texttt{argmax}_{i,s,b}\big{\{}\mathbb{E}\left[\mathsf{D}_{\mathrm{KL}}\big{(}\pi_{\ell}(\bm{\theta}|\mathbb{D}_{\ell})\,\big{|}\!\big{|}\,\pi_{\ell}(\bm{\theta})\big{)}\right]\big{\}},$ (19)
where $\pi_{\ell}(\bm{\theta}|\mathbb{D}_{\ell})$ is the posterior probability in Eq. (16) corresponding to the measurement results one would obtain by measuring at a time $t_{i}$, initial state $s$ and measurement basis $b$, and enlarging the data set as $\mathbb{D}_{\ell^{\prime}}=\mathbb{D}_{\ell}\cup\delta\mathbb{D}_{\ell^{\prime}}$. In the expression above, we are making use of the Kullback-Leibler divergence of Eq. (10) between the posterior and the prior, searching for a measurement configuration that maximizes the relative entropy between the prior and any of the possible posteriors, such that one gains the maximum amount of information at each Bayesian step. Therefore, not only the data set is enlarged sequentially $\mathbb{D}_{\ell}\supset\mathbb{D}_{\ell-1}$, but also the specific times are chosen adaptively. In light of the fixed set of measurement times in the frequentist approach $\{t_{i}:\,i\in\mathbb{I}_{t}\}$, we note that the total set of updated times after $N_{\rm B}$ Bayesian steps $\{t_{\ell},\ell\in\{0,\cdots,N_{\rm B}\}\}$ can be very different, and that is the reason why we use a different notation. Computing Eq. (19) can be quite time-consuming, especially when dealing with a large number of parameters. In practice, long computation times may lead to a reduction in the frequency of experimental shots, which is undesirable. To avoid this, we can take tens or hundreds of shots at each step before computing again the optimal measurement configuration. This will not change the results significantly, since a Bayesian update of a single shot does not change the prior much and the optimal measurement time of next step remains very similar to the previous one.
We note that the Bayesian approach is ultimately related to the maximum-likelihood estimation in the asymptotic limit $N_{\rm shot}\rightarrow\infty$. When the variance of the prior is small, the maxima of the Kullback-Leibler divergence in Eq. (10) between prior and posterior are localized at the optimal measurement configurations obtained by minimizing the asymptotic covariance matrix of the maximum likelihood-estimation in Eq. (13). Additionally, if we make a single Bayesian update in which we take a very big number of shots at several measurement configurations, the likelihood function relating the posterior and the prior will be a joint distribution of independent binomials. Given the big number of shots taken in this single step, the posterior distribution will be mainly shaped by the likelihood function, which contains most of the information about the parameters. The position of the maximum of this PDF will be located at the most likely value of the parameters, and therefore it is in agreement with the maximum-likelihood estimator that takes measurements at those measurement configurations. The Bayesian approach has the advantage that, at each step, we measure at the most convenient configuration that maximizes the expected information gain, and this can lead to an overall reduction of the number of measurements needed. Also, as noted above, the final estimate relies on a probability distribution, so that we can immediately derive confidence intervals without requiring any asymptotic limit.
For the Bayesian approach, where we use sequential Monte Carlo [112] to implement it numerically, the computational complexity is dominated by the computation of the best measurement configuration. At each step in the Bayesian approach, we need to compute the expected posterior of the measurement configurations, and then select the measurement configuration maximizing the Kullback-Leibler divergence between posterior and prior. The complexity of this calculation depends on the number of particles in the particle filter, $N_{\rm filter}$, the number of initial states, the number of measurement bases, the number of measurement outcomes and the number of measurement times considered. Therefore, we have a complexity $O(N_{\rm filter}\times|\mathbb{S}_{0}|\times|\mathbb{M}_{b}|\times|\mathbb{M}_{m_{b}}|\times|\mathbb{I}_{t}|)$. At each one of these $N_{\rm filter}\times|\mathbb{S}_{0}|\times|\mathbb{M}_{b}|\times|\mathbb{M}_{m_{b}}|\times|\mathbb{I}_{t}|$ operations we need to evaluate the likelihood function $p^{\rm TL}_{\ell}(\delta\mathbb{D}_{\ell}|\bm{\theta})$ to compute the posterior weight. Therefore if the posterior function is costly to compute we could think of precomputing it on a grid and then make linear interpolation to reduce the computational cost, as the authors of Ref. [113] do. $N_{\rm filter}$ may need to increase exponentially with the number of parameters, which makes the frequentist approach more suitable in the case of large number of parameters. However, this can be moderated by the specific structure of the model and how concentrated the posterior is.
Lindblad-like quantum tomography for non-Markovian dephasing
@@
Figure 1: Scheme for $\ell$QT protocol of a dynamical dephasing map: In the case of pure dephasing, the $\ell$QT scheme reduces to a problem of Ramsey estimation. In the QIP, the qubit is initialized in a single state in the equator of Bloch’s sphere $\rho_{0,0}=\outerproduct{+}{+}$ | by a resonant driving inducing a $\pi/2$ pulse on a reference state $\rho_{0}=\outerproduct{0}{0}$ |. The qubit then evolves for different times $t_{i}:i\in\mathbb{I}_{t}$, after which it is subjected to another resonant $\pi/2$ pulse, which effectively rotates to the Pauli measurement basis $x$. As a consequence of the device dephasing noise with actual noise $\bm{\theta}_{\star}$ parameters, the $m_{x}\in\{0,1\}$ outcomes are collected as relative frequencies $\{f_{+,i,x}(m_{x}|\bm{\theta}_{\star})\}$ which will decay in time. This data can be collected by distributing the measurement shots among pre-fixed evolution times, connecting to the frequentist estimation protocol. Alternatively, in a Bayesian approach, one collects the data by successive $\ell$-update steps $\delta\mathbb{D}_{\ell}=\{N_{+,i_{\ell},x,m_{x,\ell}}(\bm{\theta}_{\star})\}$ In the lower half, we show the classical part of the estimation, which starts by solving the time-local master equation for a particular parametrized power spectral density $S_{\bm{\theta}}(\omega)$ and filter function. This leads to the theoretically predicted probabilities $\{p^{\rm TL}_{+,i}(m_{x}|\bm{\theta})\}$. These probabilities and the relative frequencies are either fed into a log-likelihood cost function $\mathcal{C}_{\rm TL}(\bm{\theta})$ that must be minimized in order to obtain the frequentist estimate $\bm{\theta}_{\rm F}$, or used in a Bayesian step to update our prior knowledge of the noise parameters $\pi_{\ell}(\bm{\theta}|\mathbb{D}_{\ell})$, converging to the final Bayesian estimate $\bm{\theta}_{\rm B}$. In the frequentist approach, by calculating the covariance matrix, we can estimate the precision and find the optimal measurement times, which will depend on how many noise parameters we aim at estimating, and changes with their specific values. In the Bayesian approach, knowledge about the noise parameters is updated sequentially by selecting evolution times that are expected to yield the greatest information gain at each step, thereby implementing feedback and adaptive data acquisition from the quantum device.
@@
Let us formulate the $\ell$QT for the time-local master equation of Eq. (4) for the dephasing of a single qubit, including temporal correlations. As discussed in the Methods section, either in a semi-classical or quantum-mechanical model of pure dephasing, the qubit dynamics can be described by the time-local master equation in Eq. (4) with a simple Hamiltonian $H(t)=\textstyle\frac{1}{2}\omega_{0}\sigma_{z}$ and a single jump operator $L(t)=\sigma_{z}$, both of which are time independent
@@
In a semi-classical model, $C(t,t^{\prime})=\mathbb{E}[\delta\tilde{\omega}(t)\delta\tilde{\omega}(t^{\prime})]$ is the auto-correlation function of a stochastic process $\delta\tilde{\omega}(t)$ representing frequency fluctuations. This master equation is the result of averaging over the stochastic process $\rho(t)=\mathbb{E}[\tilde{\rho}(t)]$, and is valid for a random process in a second-order cumulant expansion known as the fast-fluctuation expansion or, alternatively, for a Gaussian random process with arbitrary correlation times $\tau_{\rm c}$, as discussed in the Methods section. Alternatively, for a fully quantum-mechanical dephasing model, $C(t,t^{\prime})=\textstyle\frac{1}{2}{\rm Tr}_{B}\{B_{I}(t)B_{I}(t^{\prime})\rho^{\rm ss}_{\rm B}\}$ is the auto-correlation function on the stationary state of the environment/bath $\rho^{\rm ss}_{\rm B}$, which induces fluctuations in the qubit frequency via the bath operators $B_{I}(t)$. In this case, the time-local master equation is the result of tracing over the bath degrees of freedom $\rho(t)={\rm Tr}_{B}\{\rho_{\rm SB}(t)\}$, and is valid in a second-order cumulant expansion also discussed in the Methods section. Assuming wide-sense stationarity, i.e., the mean of the stochastic process is constant and the auto-correlation function only depends on $t^{\prime}-t$, we can introduce the power spectral density (PSD) of the noise
@@
It will be useful to define the symmetrized auto-correlation function and the symmetrized PSD as $\bar{C}(t,t^{\prime})=\frac{1}{2}(C(t,t^{\prime})+C(t^{\prime},t))$ and $\bar{S}(\omega)=\frac{1}{2}(S(\omega)+S(-\omega))$, since for dephasing noise only the symmetric part of the auto-correlation function and the symmetric part of the PSD influence the time evolution, as shown in the Methods section.
@@
As discussed in the previous sections, we focus on an effective parametrization of the Hamiltonian and dissipation matrix that reduces the search space. Considering the expression in Eq. (23), the dephasing rate will depend on a certain number $n$ of real-valued noise parameters $\bm{\theta}_{\!\star}\in\Theta=\mathbb{R}^{n}$ via the PSD $S_{\bm{\theta}_{\!\star}}\!(\omega)$, where we have made explicit its parametrization. Since the Hamiltonian of the qubit is very simple, and only depends on the transition frequency $\omega_{0}$ that is known with a high precision using spectroscopic methods, it need not be included in the learning. Technically, this means that we can assume that the driving used to initialize and measure the qubit is resonant with the transition (see Fig. 1), and work in the rotating frame presented in the Methods section. Likewise, we have only one possible jump operator $L=\sigma_{z}$, such that the learning can focus directly on the estimation $\bm{\hat{\theta}}$ of the actual noise parameters $\bm{{\theta}}_{\star}$, which translate into the estimation of the dephasing rate $\hat{\gamma}(t)$ via Eq. (23). Therefore, $\ell$QT becomes a non-linear minimization problem for the cost function in Eq. (5). Before giving more details on this problem, we discuss relevant properties of the dephasing quantum dynamical map.
@@
For a white-noise noise model with a vanishing correlation time $\tau_{\rm c}=0$, one has a flat PSD $S(\omega)=c$ and a constant dephasing rate $\gamma(t)=c/2$. This leads to an exponential decay of the coherences $p_{+,i,x}(m_{x})=\textstyle\frac{1}{2}\big{(}1+(-1)^{m_{x}}\mathord{\rm e}^{-t_{i}/T_{2}}\big{)}$, where $m_{x}\in\{0,1\}$ correspond to the projective measurements on $\ket{+},\ket{-}$, respectively, and we have introduced a decoherence time $T_{2}=2/c$. If the real system is affected by white dephasing noise, there is thus a single noise parameter to learn $\theta_{\star}=c_{\star}$ or, alternatively, the real decoherence time $\theta_{\star}=T_{2\star}$. We note that this procedure is in complete agreement with the LQT based on the corresponding Lindblad master equation in Eq. (2). On the other hand, for a time-correlated dephasing noise with a structured PSD, the coherence decay will generally differ from the above exponential law, with the exception of the long-time regime $t_{i}\gg\tau_{\rm c}$, where $p_{+,i,x}(m_{x})\approx\textstyle\frac{1}{2}\big{(}1+(-1)^{m_{x}}\mathord{\rm e}^{-t_{i}/T_{2}}\big{)}$ and one finds an effective decoherence time controlled by the static part of the PSD $T_{2}=2/S(0)$. As $t_{i}$ increases towards $\tau_{\rm c}$, the decay will no longer be a time-homogeneous exponential, which can actually be a consequence of (but not a prerequisite for) a non-Markovian quantum evolution. In this more general situation, we will have more noise parameters $\bm{\theta}_{\star}$ to learn.
@@
is non-Markovian when it is not CP-divisible. In Eq. (30), we have introduced the following time-dependent probability for the occurrence of phase-flip errors